OpenStudy (anonymous):

solve by graphing... x+y=-1 2x-y=4 show stepps plaes

OpenStudy (anonymous):

Let x = -1 - y so 2x = 2(-1 - y) -2 -2y - y = 4 -3y = 6 y = -2 and substitute this into the first equation and you can find the x which is x = -1 - (-2) = 1